OpenStudy (anonymous):

ANSWER FOR MEDAL + FAN Ari’s chickens eat their pellets at a constant rate throughout each day. What is the slope of the line that represents the ounces of pellets eaten? Ari’s Chicken Pellets Days 2 4 6 8 Ounces of Pellets remaining 39 33 27 21

OpenStudy (anonymous):

@Preetha @iGreen (i LIKE YOU BECAUSE YO HELP ME THROUGH THE QUESTION)

OpenStudy (igreen):

Just take any two points from the table. Let's take (2, 39) and (4, 33) Now let's plug them into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{33-39}{4-2}\) Can you simplify that? @im.celibate

OpenStudy (anonymous):

-3 OVER 2...

OpenStudy (igreen):

No, check again.

OpenStudy (igreen):

33 - 39 isn't -3..

OpenStudy (anonymous):

Oh..Yeah.. -6

OpenStudy (texaschic101):

reduce it

OpenStudy (anonymous):

SO would the answer be -3 because -6 dived by 2 is -3...

OpenStudy (igreen):

Yep, that gives you -6 / 2, which is?

OpenStudy (anonymous):

-3, thank you so much @iGreen

OpenStudy (igreen):

Yep, you got it.